OpenStudy (anonymous):

Identify the Exponent Law shown in the example below. (a2)3 = a2·3 = a6

OpenStudy (unklerhaukus):

\[\left(a^n\right)^m=a^{n\cdot m}\]\[ (a^2)^3 = a^{2·3} = a^6\]